# KNN-MMD: Cross Domain Wireless Sensing via Local Distribution Alignment

### v2

@@ III-D1 Preliminary Classification
After performing dimension reduction, we utilize the KNN algorithm to classify the testing set based on the support set. For more detailed information on the KNN algorithm, please refer to the paper [21]. Briefly, for each sample in the testing set, we identify the nearest $k$ samples from the support set using Gaussian distance as the distance measure. Subsequently, we determine the most frequent label among the identified neighbors as the classification result. In the case of a tie, we select the label of the nearest neighbor as the classification result.
Next, we select the top $p\%$ samples from the testing set to construct the help set based on the confidence level of classification. In our approach, the confidence is measured by calculating the distance between each sample in the testing set and its nearest neighbor in the support set with the same label. A smaller distance indicates higher confidence. Through experimentation, we have found this method to be effective. For instance, in the four tasks conducted in our experiment, when setting $p\%$ to $50\%$, the classification accuracy of the selected top $p\%$ samples consistently exceeds 90%, regardless of the values of $k$ (the number of neighbors in KNN) and $n$ (the number of samples per category in the support set for n-shot tasks). Hence, we consider these samples as the help set for the subsequent steps.
@@ III-D2 Training Network with Local Alignment
$L=L_{cls}+L_{MMD}^{local}+L_{MMD}^{global}\ ,$ (10)
Firstly, $L_{cls}$ represents the traditional cross-entropy loss function, which utilizes data solely from the training set. It is used to train the model to learn the basic classification capacity. Secondly, $L_{MMD}^{local}$ represents the local MMD, which calculates the MK-MMD between the help set and the training set within each category, as shown in Eq. (11):
@@
where $E$ is the embedding result of Resnet, $K$ is the given kernal list, and $Y$ is the sample label. It’s used to realize our local alignment. Additionally, we add an extra global MMD, which computes the MK-MMD between the entire help set and training set, following the conventional MMD approach, as represented in Eq. (12):
@@
As mentioned previously, our primary focus lies on the local MMD rather than the global MMD. However, we include the global MMD to address potential challenges arising from an imbalanced distribution of samples among different labels within the help set. By incorporating the global MMD, we can mitigate this issue to some extent. For example, if we ensure $P_{s}(x|y)=P_{t}(x|y)$ for all $y$ except $y=i$, and also make $P_{t}(x)=P_{s}(x)$, we can achieve $P_{s}(x|y=i)=P_{t}(x|y=i)$. Nonetheless, it is important to note that this approach has limitations when there is a significant lack of certain categories. Alternatively, we can employ other methods to generate the help set and address category imbalance. For example, within the testing set, we can directly select the top $p\%$ samples within each category or choose a fixed number of samples with the highest confidence level within each category. However, this may impact the accuracy of the help set to some extent.
@@
Algorithm 1 Training Method
@@
    $[X^{train},Y^{train}]$: Training Set
    $[X^{support},Y^{support}]$: Support Set
    $[X^{help},Y^{help}]$: Help Set
    $\theta$: Network Parameter
    $K$: Kernal List of MK-MMD
    $e^{min}$: Minimum Training Epoch
    $e^{max}$: Maximum Training Epoch
    $e^{threshold},\alpha\geq 1,\beta\leq 1$: Hyper-parameters for Early Stop
1:Initialize network Resnet and MLP
2:Set best history loss $loss^{best}:=\infty$
3:Set best history accuracy $acc^{best}:=0$
4:Set epoch counters $e^{acc}:=1$, $e^{loss}:=1$
5:for $i=1$ to $e^{max}$ do
6:     Calculate embedding result $E:=\text{Resnet}(X)$
7:     Calculate classification result $\hat{Y}=\text{MLP}(E)$
8:     Calculate loss function $l^{train}$ according to Eq. 10
9:     Update network parameter $\theta:=\theta-\frac{dl^{train}}{d\theta}$
10:     Calculate cross entropy loss $l$ in help set
11:     Calculate classification accuracy $a$ in help set
12:     if $a\geq acc^{best}$ then
13:         Update best history acc $acc^{best}:=a$
14:         Reset counter $e^{acc}:=1$
15:         Save current parameter $\theta$
16:     else
17:         Increment counter $e^{acc}:=e^{acc}+1$
18:     end if
19:     if $l\leq loss^{best}$ then
20:         Update best history loss $loss^{best}:=l$
21:         Reset counter $e^{loss}:=1$
22:         Save current parameter $\theta$
23:     else
24:         Increment counter $e^{loss}:=e^{loss}+1$
25:     end if
26:     if $i==e^{min}$ then
27:         Increase best history loss $loss^{best}:=loss^{best}\cdot\alpha$
28:         Decrease best history acc $acc^{best}:=acc^{best}\cdot\beta$
29:         Reset counters $e^{acc}:=1$, $e^{loss}:=1$
30:     end if
31:     if $i>e^{min}\ \&\&\ e^{loss}>e^{threshold}\ \&\&\ e^{acc}>e^{threshold}$ then $\triangleright$ Early Stop
32:         Break
33:     end if
34:end for
@@ IV-A1 WiGesture Dataset
The WiGesture Dataset is a gesture recognition and people identification dataset. We use the dynamic part of the dataset that include 6 action id and 8 people id. The dataset is collected by 8 college students in a meeting room. The ESP32-S3 serves as receiver and a home Wi-Fi router serves as transmitter. The six actions include left-right, forward-backward, and up-down motions, clockwise circling, clapping, and waving. In the gesture recognition task, we consider different people as distinct domains. Similarly, in the people identification task, we treat different actions as separate domains.
@@ IV-A2 WiFall Dataset
The data collection scenario and device are the same as the WiGesture dataset, as shown in Fig. 7. In the dataset, we have a total of 10 volunteers with an average age of 23.17, height of 1.75m, and BMI of 23.76. The activities include walking, jumping, sitting, standing up, and falling, which also consists of many types like forward fall, left fall, right fall, seated fall, and backward fall, shown as Fig. 8. Similar to the WiGesture Dataset, we consider different people as distinct domains. In our experiment, we use WiFall dataset to conduct action recognition tasks, where we identify five different actions, as well as fall detection tasks, where we only distinguish between fall and other actions.
@@ IV-B Experiment Setup
Task Training Set Support Set Testing Set
Gesture Recognition People ID 1-7 $n$ samples for each gesture in People ID 0 samples excluding support set in People ID 0
People Identification Action ID 1-5 $n$ samples for each person in Action ID 0 samples excluding support set in Action ID 0
Fall Detection & Action Recognition People ID 1-9 $n$ samples for each action in People ID 0 samples excluding support set in People ID 0
@@ IV-C Experiment Result
Method Scenario Gesture Recognition People Identification Fall Detection Action Recognition Average Accuracy Resnet18[20] in-domain 80.75% 86.75% 91.88% 70.50% 82.47% zero-shot 40.84% 70.50% 59.86% 26.00% 49.30% Siamese[23] one-shot 70.40% 82.87% 60.62% 38.95% 63.21% AutoFi (MLP-based)[43] one-shot 24.62% 24.71% 50.88% 23.59% 30.95% AutoFi (CNN-based)[43] one-shot 27.05% 36.14% 48.05% 26.95% 34.55% Yang et al.[16] one-shot 67.21% 74.22% 59.75% 48.52% 62.43% Ding et al.[44] one-shot 39.14% 70.94% 61.56% 30.37% 50.50% CrossFi[18] one-shot 91.72% 93.01% 80.93% 49.62% 78.82% KNN [21] one-shot 83.02% 82.67% 49.63% 46.87% 65.55% KNN-MMD (Ours) one-shot 93.26% 81.84% 77.62% 75.30% 82.01% Ablation Study one-shot 69.87% 73.78% 84.03% 74.06% 75.44% MMD[25] zero-shot 47.92% 67.25% 74.32% 45.61% 58.75% MK-MMD[39] zero-shot 40.36% 66.47% 72.26% 43.72% 55.70% DANN[28] zero-shot 41.41% 67.18% 74.06% 35.99% 54.66% ADDA[45] zero-shot 42.71% 65.43% 62.81% 36.08% 51.76% GFK+KNN[26] zero-shot 30.79% 51.50% 53.72% 34.17% 42.55% CrossFi[18] zero-shot 64.81% 72.79% 74.38% 40.46% 63.11% Tian et al.[36] zero-shot 68.13% 55.86% 61.72% 42.10% 56.95% EEG[37] zero-shot 59.75% 64.63% 69.53% 42.15% 59.02%
@@ IV-D3 DAL Method
As mentioned before, the global alignment operation of traditional DAL methods cannot guarantee the model performance. In Fig. 11, we illustrate the UMAP result of different methods. It can be seen that the performance of the global alignment method MK-MMD does not differ much from ResNet, which even has no alignment operation. Even though they can split samples in the source domain (training set) well, achieving an accuracy over 99%, they cluster samples in the target domain (testing set) chaotically. In contrast, our KNN-MMD still performs well in the target domain. As shown in Fig. 11c, it successfully aligns samples from the source domain and target domain within each category. Furthermore, in Fig. 11d, we show the upper bound performance of KNN-MMD, where we provide the ground truth label of the help set instead of pseudo-labels. We notice that the cluster performance and accuracy of KNN-MMD and the upper bound do not have a significant difference, which proves the efficiency of our help set construction method. Finally, even though our early stop method can be used in most DAL methods, as the support set does not directly participate in training, limited by the model principle, even if we make it stop at the point with the best testing accuracy, it still has a poor performance, as shown in Fig. 6.

### v3

@@ III-D1 Preliminary Classification
After performing dimension reduction, we utilize the KNN algorithm to classify the testing set based on the support set. For more detailed information on the KNN algorithm, please refer to the paper [21]. Briefly, for each sample in the testing set, we identify the nearest $k$ samples from the support set using Euclidean distance as the distance measure. Subsequently, we determine the most frequent label among the identified neighbors as the classification result. In the case of a tie, we select the label of the nearest neighbor as the classification result.
Next, we select the samples from the testing set that have the top $p\%$ confidence level of classification to construct the help set. In our approach, confidence is measured by calculating the distance between each sample in the testing set and its nearest neighbor in the support set with the same label. A smaller distance indicates a higher confidence level. The detailed construction process is outlined in Algorithm 1.
Algorithm 1 Help Set Construction
Require:
    $[X^{support},Y^{support}]$: Support Set
    $[X^{test}]$: Testing Set
    $k$: Neighbor Amount in KNN
    $p\%$: Use Top-$p\%$ Confidence Testing Samples as Help Set
    $\epsilon$: Small constant for numerical stability
[…]
2:for each sample $x_{i}\in X^{test}$ do
3:     $N\leftarrow\text{KNN}(x_{i},X^{support},k)$triangleright Find $k$ nearest neighbors in support set
4:     $\hat{y}_{i}\leftarrow\text{majorityVote}(Y^{support}_{N})$triangleright Predict pseudo-label
5:     $X^{support}_{\hat{y}_{i}}\leftarrow\{x_{j}\in X^{support}\mid Y^{support}_{j}=\hat{y}_{i}\}$
6:     $x^{*}\leftarrow\text{argmin}_{x\in X^{support}_{\hat{y}_{i}}}\text{distance}(x_{i},x)$
7:     $d_{i}\leftarrow\text{distance}(x_{i},x^{*})$
8:     $c_{i}\leftarrow\frac{1}{d_{i}+\epsilon}$triangleright Compute confidence score
[…]
10:end for
[…]
12:$n\leftarrow\lfloor p\%\times|X^{test}|\rfloor$triangleright Number of samples to select
[…]
Through experimentation, we have found this method to be effective. For instance, in the four tasks conducted in our experiment, when setting $p\%$ to $50\%$, the classification accuracy of the selected top $p\%$ samples consistently exceeds 90%, regardless of the values of $k$ (the number of neighbors in KNN) and $n$ (the number of samples per category in the support set for n-shot tasks). Hence, we consider these samples as the help set for the subsequent steps.
@@ III-D2 Training Network with Local Alignment
$L=L_{cls}+\alpha_{1}L_{MMD}^{local}+\alpha_{2}L_{MMD}^{global}\ ,$ (11)
where $\alpha_{1},\alpha_{2}$ represent the weights. In the loss function, firstly, $L_{cls}$ represents the traditional cross-entropy loss function, which utilizes data solely from the training set. It is used to train the model to learn the basic classification capacity. Secondly, $L_{MMD}^{local}$ represents the local MMD, which calculates the MK-MMD between the help set and the training set within each category, as shown in Eq. (12):
@@
where $M$ represents the total category amounts, $E$ is the embedding result of ResNet, $K$ is the given kernel list, and $Y$ is the sample label. It’s used to realize our local alignment. Additionally, we add an extra global MMD, which computes the MK-MMD between the entire help set and training set, following the conventional MMD approach, as represented in Eq. (13):
@@
As mentioned previously, our primary focus lies on the local MMD rather than the global MMD. However, we include the global MMD to address potential challenges arising from an imbalanced distribution of samples among different labels within the help set, as our construction method does not guarantee balanced labels. Since the MMD is based on the statistical properties of data distribution, a small sample size can introduce random bias. By incorporating the global MMD, we can mitigate this issue to some extent. For example, if we ensure $P_{s}(x|y)=P_{t}(x|y)$ for all $y$ except $y=i$, and also make $P_{t}(x)=P_{s}(x)$, we can achieve $P_{s}(x|y=i)=P_{t}(x|y=i)$.
Nonetheless, it is important to note that this approach has limitations when there is a significant lack of certain categories. Alternatively, we can employ other methods to generate the help set and address category imbalance. For example, within the testing set, we can directly select the top $p\%$ samples within each category or choose a fixed number of samples with the highest confidence level within each category. However, this may impact the accuracy of the help set to some extent.
@@
Remark 2.
Here, we want to emphasize the principles and validity of our early stopping approach. During training, we only have labeled training data from the source domain, a small number of labeled support samples from the target domain, and a large amount of unlabeled testing data from the target domain. It is important to note that throughout the entire training process, we do not access any ground truth labels for the testing data. Initially, we obtain pseudo labels for the testing data using KNN by comparing their distances with the labeled support data. We then select a subset of data with the highest confidence levels as the help set, ensuring that the support set is not included in the help set.
Next, we train a neural network using the training data to enhance its classification capacity and utilize both the training data and help data for local alignment. Consequently, during the entire training phase, the network does not access any data from the support set or any ground truth labels from the testing set. Since the support set shares the same distribution as the testing set (both originate from the target domain), we can use the model’s performance on the support set as a signal for early stopping. Although the help set is selected using support data, which may introduce some bias, we have found this method to be effective through experimentation. Furthermore, since the ground truth for the testing set remains unavailable throughout the training and validation phases, our method does not risk any potential information leakage.
Algorithm 2 Training with Adaptive Early Stopping
@@
    $[X^{train},Y^{train}]$: Training data and labels
    $[X^{support},Y^{support}]$: Support set for domain adaptation
    $[X^{help},Y^{help}]$: Help set for validation
    $\theta$: Network parameters
    $K$: Kernel list for MK-MMD
    $e^{min}$: Minimum training epochs
    $e^{max}$: Maximum training epochs
    $e^{threshold}$: Patience threshold for early stopping
    $\alpha\geq 1$: Loss relaxation factor
    $\beta\leq 1$: Accuracy relaxation factor
1:Initialize:
[…]
3:$loss^{best}\leftarrow\infty$, $acc^{best}\leftarrow 0$, $\triangleright$ Initialize best validation loss and best validation accuracy
4:$e^{acc},e^{loss}\leftarrow 1$triangleright Set Epoch counters
5:$\theta^{best}\leftarrow\theta$triangleright Store best parameters
6:for epoch $i=1$ to $e^{max}$ do
7:     Forward Pass:
8:     $E\leftarrow\text{ResNet}(X^{train})$triangleright Extract features
9:     $\hat{Y}\leftarrow\text{MLP}(E)$triangleright Compute predictions
10:     Loss Computation:
11:     $l^{train}\leftarrow\mathcal{L}(\hat{Y},Y^{train})$triangleright Eq. 11
12:     $\theta\leftarrow\theta-\eta\nabla_{\theta}l^{train}$triangleright Parameter update
13:     Validation on Help Set:
14:     $l\leftarrow\text{CrossEntropy}(X^{help},Y^{help})$triangleright Validation loss
15:     $a\leftarrow\text{Accuracy}(X^{help},Y^{help})$triangleright Validation accuracy
16:     Accuracy Tracking:
17:     if $a\geq acc^{best}$ then
18:         $acc^{best}\leftarrow a$
19:         $e^{acc}\leftarrow 1$
20:         $\theta^{best}\leftarrow\theta$triangleright Save improved parameters
21:     else
22:         $e^{acc}\leftarrow e^{acc}+1$
23:     end if
24:     Loss Tracking:
25:     if $l\leq loss^{best}$ then
26:         $loss^{best}\leftarrow l$
27:         $e^{loss}\leftarrow 1$
28:         $\theta^{best}\leftarrow\theta$triangleright Save improved parameters
29:     else
30:         $e^{loss}\leftarrow e^{loss}+1$
31:     end if
32:     Relaxation at Minimum Epochs:
33:     if $i==e^{min}$ then
34:         $loss^{best}\leftarrow loss^{best}\cdot\alpha$triangleright Relax loss threshold
35:         $acc^{best}\leftarrow acc^{best}\cdot\beta$triangleright Relax accuracy threshold
36:         $e^{acc},e^{loss}\leftarrow 1$triangleright Reset counters
37:     end if
38:     Early Stopping Condition:
39:     if $i>e^{min}$ and $e^{loss}>e^{threshold}$ and $e^{acc}>e^{threshold}$ then
40:         break $\triangleright$ Stop training when both metrics stagnate
41:     end if
42:end for
43:Return $\theta^{best}$triangleright Return best parameters found
@@ IV-A1 WiGesture Dataset
The WiGesture Dataset is a gesture recognition and people identification dataset. We use the dynamic part of the dataset that includes 6 action id including left-right, forward-backward, and up-down motions, clockwise circling, clapping, and waving and 8 people id. The dataset is collected by 8 college students in a meeting room. The ESP32-S3 serves as receiver and a home Wi-Fi router serves as transmitter. The six actions include left-right, forward-backward, and up-down motions, clockwise circling, clapping, and waving. In the gesture recognition task, we consider different people as distinct domains. Similarly, in the people identification task, we treat different actions as separate domains.
@@ IV-A2 WiFall Dataset
The data collection scenario and device are the same as the WiGesture dataset, as shown in Fig. 8. In the dataset, we have a total of 10 volunteers with an average age of 23.17, height of 1.75m, and BMI of 23.76. The activities include 5 types: walking, jumping, sitting, standing up, and falling, which also consists of many types like forward fall, left fall, right fall, seated fall, and backward fall, shown as Fig. 9. Similar to the WiGesture Dataset, we consider different people as distinct domains. In our experiment, we use WiFall dataset to conduct action recognition tasks, where we identify 5 different actions, as well as fall detection tasks, where we only distinguish between fall and other actions.
@@ IV-B Experiment Setup
In the experiment, we split the dataset into segments of 1 second each. In the WiGesture dataset, each action performed by each individual lasts for 1 minute, resulting in 60 samples per action. There are 6 gestures and 8 participants in total. For the gesture recognition task, we use participants with IDs 1-7 as the training set, which includes 420 samples for each class, totaling 2,520 samples. The last 360 samples (60 samples for each class) from participant ID 0 are used as the target domain. In the people identification task, we use action IDs 1-5 for the training set, which consists of 300 samples per class, totaling 2,400 samples. The last 480 samples (60 samples for each class) from action ID 0 serve as the target domain.
In the WiFall dataset, each action performed by each individual lasts for 12 seconds, resulting in 12 samples, except for the action “fall”. For the “fall” action, each participant repeats it 48 times, with each repetition lasting 1 second, corresponding to 1 sample. There are 5 actions, including “fall”, across 10 participants. For the fall detection task, we combine all non-fall actions into a single category and use participants with IDs 1-9 as the training set, which includes 432 samples for each class, totaling 864 samples. The last 96 samples (48 samples for each class) from participant ID 0 are used as the target domain. In the action recognition task, to maintain label balance between classes, we randomly drop $\frac{3}{4}$ of the samples from the “fall” class. We then use participants with IDs 1-9 as the training set, which includes 108 samples per class, totaling 540 samples. The last 60 samples (12 samples for each class) from participant ID 0 are used as the target domain.
@@
Task Dataset Training Set Support Set (Also Used as Valid Set) Testing Set Gesture Recognition WiGesture People ID 1-7 $n$ samples for each gesture in People ID 0 samples excluding support set in People ID 0 People Identification WiGesture Action ID 1-5 $n$ samples for each person in Action ID 0 samples excluding support set in Action ID 0 Fall Detection WiFall People ID 1-9 $n$ samples for each action in People ID 0 samples excluding support set in People ID 0 Action Recognition WiFall People ID 1-9 $n$ samples for each action in People ID 0 samples excluding support set in People ID 0 Task Class Amount Training Set Size Support Set Size Testing Set Size Gesture Recognition 6 $6\times 420=2,520$ $6\times n$ $6\times(60-n)$ People Identification 8 $8\times 300=2,400$ $8\times n$ $8\times(60-n)$ Fall Detection 2 $2\times 432=864$ $2\times n$ $2\times(48-n)$ Action Recognition 5 $5\times 108=540$ $5\times n$ $5\times(12-n)$
@@
To evaluate the efficiency of our method, we compare it with several representative cross-domain methods in the fields of Wi-Fi sensing and machine learning. For fairness, we have made every effort to replicate other works according to the descriptions provided in their respective papers. For any missing settings in these papers, we adopt the same settings as our method.
Wi-Fi sensing methods can be classified into two types: signal processing methods (which extract domain-independent information) and feature learning methods [43]. Domain-independent methods typically require multiple RX-TX pairs or antennas for feature extraction (for example, WiDar3.0 [44] uses 1 TX and 6 RXs), which is not applicable in our experimental setting, where only single RX-TX pairs and a single antenna are available. For feature learning methods, most works follow a similar paradigm: they first train a feature extractor and then compare the similarity of the extracted features from the source and target domains for classification. We have selected four representative methods for comparison: Ding et al. [45], AutoFi [46], Yang et al. [47], and CrossFi [18]. Ding et al. [45] propose using a CNN trained in the source domain as a feature extractor and then use cosine similarity in the feature embedding for similarity calculations. Additionally, they employ a low-pass Butterworth filter for denoising. Based on this, AutoFi [46] is introduced (without the low-pass denoising), proposing a pre-training method to enhance the feature extractor’s capacity in the neural network. In Yang et al. [47], a different similarity calculation method is employed, specifically a Siamese network [23]. They also choose a CNN combined with Bi-LSTM as the feature extractor. Building on this, CrossFi [18] uses a ResNet-based Siamese architecture, introducing a template generation method to enhance feature representation and an attention-based similarity calculation method to better capture relationships between samples. In CrossFi, the application of the Siamese network is extended to a zero-shot scenario. In feature learning methods, the two most important factors are the feature extractor and the similarity calculation method. We selected these four methods because they represent different powerful feature extractors and similarity calculation techniques.
In the field of Wi-Fi sensing, most works are based on the feature learning paradigm mentioned above. Even some works combined with Domain Adaptation Learning (DAL) methods, such as Yang et al. [47] and CrossFi [18], which also utilize the MK-MMD loss, do not rely solely on it. To illustrate the performance of the DAL methods, we have chosen several popular methods in machine learning, including adversarial learning-based DANN [28] and ADDA [48], as well as feature-space alignment methods such as MK-MMD [40] and GFK [26]. Since these methods do not depend on specific network structures, we use the same ResNet architecture as in this paper. Additionally, we employ the same kernel functions in MK-MMD as in our KNN-MMD method. Furthermore, we also compare the methods of Tian et al. [36] and EEG [37], which also discuss local alignment methods, with detailed descriptions provided in Section II-C. All these DAL methods fall under the zero-shot category, meaning they only require a labeled training set from the source domain and an unlabeled testing set, without the necessity for a labeled support set.
@@ IV-C Experiment Result
Method Scenario Gesture Recognition People Identification Fall Detection Action Recognition Average Accuracy ResNet18[20] in-domain 80.75% 86.75% 91.88% 70.50% 82.47% zero-shot 40.84% 70.50% 59.86% 26.00% 49.30% Siamese[23] one-shot 70.40% 82.87% 60.62% 38.95% 63.21% AutoFi (MLP-based)[46] one-shot 24.62% 24.71% 50.88% 23.59% 30.95% AutoFi (CNN-based)[46] one-shot 27.05% 36.14% 48.05% 26.95% 34.55% Yang et al.[16] one-shot 67.21% 74.22% 59.75% 48.52% 62.43% Ding et al.[49] one-shot 39.14% 70.94% 61.56% 30.37% 50.50% CrossFi[18] one-shot 91.72% 93.01% 80.93% 49.62% 78.82% KNN [21] one-shot 83.02% 82.67% 49.63% 46.87% 65.55% KNN-MMD (Ours) one-shot 93.26% 81.84% 77.62% 75.30% 82.01% Ablation Study one-shot 69.87% 73.78% 84.03% 74.06% 75.44% MK-MMD[40] zero-shot 40.36% 66.47% 72.26% 43.72% 55.70% DANN[28] zero-shot 41.41% 67.18% 74.06% 35.99% 54.66% ADDA[48] zero-shot 42.71% 65.43% 62.81% 36.08% 51.76% GFK+KNN[26] zero-shot 30.79% 51.50% 53.72% 34.17% 42.55% CrossFi[18] zero-shot 64.81% 72.79% 74.38% 40.46% 63.11% Tian et al.[36] zero-shot 68.13% 55.86% 61.72% 42.10% 56.95% EEG[37] zero-shot 59.75% 64.63% 69.53% 42.15% 59.02%
@@ IV-D3 DAL Method
As mentioned before, the global alignment operation of traditional DAL methods cannot guarantee the model performance. In Fig. 12, we illustrate the UMAP result of different methods, including:
  • $\bullet$
    ResNet (no alignment): We directly train the ResNet18 on the source domain and use it to embed samples from the target domain.
  • $\bullet$
    MK-MMD (global alignment): Building on the ResNet (no alignment), we incorporate the external MK-MMD loss function to achieve global alignment between the source and target domains.
  • $\bullet$
    KNN-MMD (shot number=1): This method, proposed in this paper, aligns the source and target domains locally (within each category) by utilizing source domain data with real labels and help set data with pseudo labels.
  • $\bullet$
    Theoretical Upper Bound (local alignment): This method, based on KNN-MMD, uses the ground truth labels from the target domain to achieve local alignment. It represents the upper bound for our KNN-MMD approach, which can be realized when all pseudo labels are accurate.
It can be seen that the performance of the global alignment method MK-MMD does not differ much from ResNet, which even has no alignment operation. Even though they can split samples in the source domain (training set) well, achieving an accuracy over 99%, they cluster samples in the target domain (testing set) chaotically. In contrast, our KNN-MMD still performs well in the target domain. As shown in Fig. 12c, it successfully aligns samples from the source domain and target domain within each category. Furthermore, in Fig. 12d, we show the upper bound performance of KNN-MMD, where we provide the ground truth label of the help set instead of pseudo-labels. We notice that the cluster performance and accuracy of KNN-MMD and the upper bound do not have a significant difference, which proves the efficiency of our help set construction method. Finally, even though our early stop method can be used in most DAL methods, as the support set does not directly participate in training, limited by the model principle, even if we make it stop at the point with the best testing accuracy, it still has a poor performance, as shown in Fig. 9.